Skip to end of metadata
Go to start of metadata

Weather:   windy and rainy this morning. May clear up in the afternoon.

Sensor Status Overview
   A number of sonics, ~7, were down this morning. It appears rain was the reason. Most are back up.
  Three Licors were down but I was able to get running again via PIO commands.
  TRH 10m and 60m on rsw06 were down due to fan current going to high. I was able to get them back
       up via PIO. The max current cutoff has been increased, will monitor.

 

 

  • No labels

4 Comments

  1. It looks like the GPS at V03 isn't seeing satellites.  Is this a known issue?  (I don't see it in TODOs.)  Since the GPS is hard-wired to the Pi-hat board, we can't power cycle it separate from the Pi.

    1. So I just rebooted it to see if that would help, but it didn't.  For the sake of at least keeping the system time in sync, I added v04 and tnw10 as servers in /etc/chrony/chrony.conf.

  2. Visited v03 this morning.  As before, GPS was sending essentially blank messages.  Reset antenna microdot connector, waited one minute, but no change.  Cycled power to entire DSM.  As soon as I was able to log back in (through Bluetooth console), GPS was alive and seeing 8 satellites.  Thought I was done, but....

    Just after making this entry, logged back in and found that the GPS was dead again.

    This behavior makes me think that a reset command also could have fixed it, but the only one I see is CFG-RST, which is a binary command.  I don't know how to send this binary command to the ublox.  Also, I've verified that the GPS power is hard-wired to Vcc on the PiHat.  If we had wired it to a DIO pin, we could reset power remotely.  (Though direct connection to a DIO pin likely would not have enough power.) 

  3. So...revisited site.  Replaced PiHat.  Note that the label on this Pi is now 59, rather than 47 (but the box is still 247).  It took about 3 min for the new ublox to get a position, but now seems fine.  Thus, something is intermittently wrong with the ublox on PiHat #47.  I'll put it in box 259 for now.